How do I solve these problems? (1) 7 3\4 + 17.85 + 6 1\2 ...(2) 9\6=x\8...(3) 1\2 +(2\3÷3\4)-(4\5×5\6)
Answer by rfer(16322)

7.75+17.85+6.5=32.1=32 1/10 either way
----------
9/6=x/8 work like a proportion
6x=72
x=12
---------------
1/2+(2/3)/(3/4)-(4/5*5/6)
1/2+2/3*4/3-4/5*5/6
1/2+8/9-20/30
135/270+240/270-225/270=600/270=2 2/9